40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 18:03:47 2014, ericq, Update, Green Locking, uPDH Box Checkup uPDH_X.JPGuPDH_Y.JPGupdhTFs.pdfupdhNoises.pdfuPDH_Aug27.zip
    Reply  Thu Aug 28 17:34:20 2014, ericq, Update, Green Locking, uPDH Box Checkup TEK00006.PNG
       Reply  Fri Aug 29 17:28:07 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Fri Aug 29 23:53:35 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Tue Sep 2 20:02:06 2014, ericq, Update, Green Locking, uPDH Box Checkup 
       Reply  Wed Sep 3 00:17:22 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Wed Sep 3 20:38:40 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Fri Sep 5 04:07:44 2014, ericq, Update, Green Locking, X end uPDH Box Replaced 
                Reply  Fri Sep 5 16:13:00 2014, ericq, Update, Green Locking, new X Green PDH measurements 6x
                   Reply  Sat Sep 6 01:48:54 2014, rana, Update, Green Locking, new X Green PDH measurements 
                      Reply  Sat Sep 6 14:49:12 2014, ericq, Update, Green Locking, new X Green PDH measurements xGComparison.pdfxGspectra.pdf
                         Reply  Sun Sep 7 17:06:38 2014, rana, Update, Green Locking, new X Green PDH measurements 
                         Reply  Mon Sep 8 14:14:13 2014, Jenne, Update, Green Locking, new Y Green PDH measurements 6x
Message ID: 10437     Entry time: Thu Aug 28 17:34:20 2014     In reply to: 10433     Reply to this: 10438   10443
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: uPDH Box Checkup 

I had noticed in the past, that the digital control signal monitor for the X end would saturate well before the ADC should saturate (C1:ALS-X_SLOW_SERVO_IN1, which is from the "output mon" BNC on the box). It turns out that there is some odd saturation happening inside the box itself.

In this scope trace, the servo input is being driven with a 0.02Vpp, 0.1Hz sine wave, gain knob at 1.0. This is bad. 

TEK00006.PNG

Evan and I poked around the board, and discover that for some reason currently unknown to us, the variable gain amplifier (AD8336) can't reach its negative rail, despite the +-12V arriving safely at its power supply pins. 

I also realized that the LF356 in the integrator stage in this box had been replaced with a LT1792 by Kiwamu in ELOG 4373. I've updated my schematic, and will upload both boxes' schematics to the DCC page Jenne created for them. (D1400293 and D1400294)

ELOG V3.1.3-